Status Report

NASA Space Station On-orbit Status 13 Aug 2003

By SpaceRef Editor
August 13, 2003
Filed under , , ,
NASA Space Station On-orbit Status 13 Aug 2003

All ISS systems continue to function nominally, except as noted
previously or below.

The crew had another very busy day, with task ranging across the board
from fluid transfer to medical procedures, physics of tools,
cinematography, cable sleuthing and battery charging, performing what
Flight Control called "a fabulous job".

CDR Yuri Malenchenko, assisted by crewmate Lu, completed another session
of the MO-5 MedOps protocol of cardiovascular evaluation during graded
exercises on the VELO cycle ergometer.  Last time done: 7/17. 
[As before, Malenchenko used the Gamma-1 ECG equipment with biomed
harness, skin electrodes, and a blood pressure and rheoplethysmograph
cuff wired to the cycle ergometer’s instrumentation panels.  For the
graded exercise, he worked the pedals after a prescribed program at load
settings of 125, 150, and 175 watts for three minutes each.  All
measurements were recorded and telemetered during Daily Orbit 2 to MCC-M,
from where the workout was monitored by a specialist.]

Science Officer Ed Lu set up and then took the periodic O-OHA
(on-orbit hearing assessment) test, a NASA EHS (environmental health
systems) examination to assess the efficacy of acoustic
countermeasures.  Yuri Malenchenko followed suit.  [The
O-OHA test involves minimum audibility measurements for each ear over a
wide range of frequencies and sound pressure levels, with the crewmembers
using individual-specific Prophonics earphones, Bose ANC headsets and the
SLM (sound level meter).  To conduct the testing, the experimenter
is supported by special ìEarQî software on the MEC (medical equipment
computer).  The baseline test is required for about Flight Day 14
for each new Expedition and is then performed once per month.  Last
time done: 7/14.]

Malenchenko continued transferring liquid waste to the empty Rodnik
water storage tanks (BV1, BV2) of 10P/Progress-247, begun
yesterday.  The operation, which was begun yesterday, today
transferred stored urine from 9 EDV-U containers into tank BV1, then BV2,
from a total of 13 filled EDVs.  The work will be finished tomorrow
(8/14).  [Yesterday’s transfer involved 4 EDVs.  In
preparation for the activity, Yuri, on 8/4, had pressurized the folded
bladders of the two spherical Rodnik water storage tanks of 10P.]

As part of regular monthly preventive maintenance of RS (Russian
segment) air ventilation systems, Ed Lu worked 1.5 hrs. in the
Funktsionalnyi-Grusovoi Blok (FGB) module for a thorough cleanup
of the VT7 and VT8 fan screens of the three SOTR (thermal control system)
gas-liquid heat exchangers (GZhT4).  The fans were powered off for
the task, later turned on again.

Dr. Lu prepared and set up the video camcorder in the U.S. Lab for
today’s scheduled EPO (Educational Payload Operations) demonstration,
postponed from last week (8/7).  [The activity demonstrated the
use of EVA and IVA (intravehicular activity) tools in space, including
IVA drill driver, EVA pistol grip tool (PGT), high-torque locker tool and
torque wrench, clamp and bracket kit, leak repair kit, body restraint
tether use, maintenance work area, tool restraints, etc.  The
resulting video will be used to supplement NASA educational materials, as
well as be featured at NASA education websites.]

After Ed Lu’s preparations of the MSG (Microgravity Science Glovebox)
yesterday for the upcoming PFMI (Pore Formation and Mobility
Investigation) payload operations, the lack of a missing cable for the
MSG proved a showstopper for today’s planned MSG/PFMI checkout. 
Without the parallel-port cable (W206), PFMI cannot run.  [The
time block allotted for today’s PFMI operations was to be used for the
"cable hunt".  If the cable was/is found, Ed Lu can
execute the PFMI equipment checkout procedure as originally scheduled for
today, which uses PFMI sample #4, connected to its electronics but not
installed in PFMI thermal chamber (TS), allowing the temperature sensors
to read nominal values and not generate errors.  The sample ampoule,
whose indium-tin-oxide coating acts as its own heater when installed in
the TS, needs to remain in the polycarbonate tube for the duration of the

In another round of regular EVA/EMU (extravehicular mobility unit)
battery maintenance, Ed Lu initiated the charge/discharge procedure on
batteries #2029 and #2030, which is expected to last ~24 hrs. 
[The helmet light and PGT (pistol grip tool) batteries, which are not
charged for preservation, were removed from the BSA (battery storage
assembly) beforehand.]

Later in the day, Ed ran the first part of a series of tests aimed at
finding out if the crew can record taped video and taped audio at the
same time.  The video-with-audio test will be downlinked at a later
time.  [As of now, the communications configuration is only able
to downlink taped video with live narration ("voice

Ed Lu conducted the daily routine maintenance of the SM’s SOZh life
support system (including ASU toilet facilities), prepared the daily IMS
inventory delta file, and completed the regular status checkup (every
Monday, Wednesday and Friday) of the autonomous Lab payload (currently
only PCG-STES010, Protein Crystal Growth-Single Locker Thermal Enclosure

Yuri performed the weekly inspection of the BRPK-1 air/water condensate

The station’s cabin atmosphere was refreshed today with gaseous oxygen
(O2) from Progress 10P O2 tanks.  This will be continued as long as
its supply lasts (originally 50 kg/110 lbs), whenever O2 partial pressure
(ppO2) decreases to 160 mmHg.  Afterwards, O2 repress will also be
performed from Progress 11P.  [The Elektron oxygen generator is
off, to conserve its service life.  The repress valve cycling, by
the crew, relies on the U.S. MCA (major constituents analyzer) or CSA-CP
(compound specific analyzer-combustion products) instruments to measure
the 10mmHg pressure difference.]

The 150 Mbps (megabytes per second) Ku-band tests performed all last
week jointly by a number of ground teams were a total success. 
[The link was reported good with no frame drop-outs for both autotrack
and open loopmodes. The error-to-noise (E/N) ratios for the high
bandwidthtests wereon the order of13.0 decibel (dB), in autotrackand open
loop mode. This compares well with the existingE/Ns at the normal 50 Mbps
data rates used regularly today. During the test, engineers were able
toincrease the downtrack error vector to about 142 kmbefore finally
experiencing some frame dropouts, greatly exceeding the 30 km
specification.  Future applications will utilize and massively
benefit from this capability.]

Yesterday and today, TsUP/Moscow conducted the periodic efficiency
test on the Service Module (SM) solar arrays, this time on the backup
channel.  Afterwards, the arrays were switched back to primary
channel.  [The efficiency tests, to monitor degradation of the
photovoltaic cells due to ultraviolet and other environmental effects,
are also conducted periodically on the USOS arrays.]

MCC-H ballistics experts continue to follow the conjunction with an
upper stage of aChinese Long March 4B (CZ-4B) vehicle launched 10/27/02
(object #27551).  Preliminary TCA (time of closest approach) remains
unchanged: Thursday (8/14), 7:41pm EDT.  The concern about this
conjunction has lessened overnight as miss distance data have remained
steady.  No maneuver is expected at this time; however, MCC-H will
continue to monitor.

Today’s CEO (crew earth observation) targets, no longer limited in the
current LVLH attitude and including the targets of the Lewis & Clark
200-year memorial locations, were Machu Picchu, Peru (nadir
Amazonian rivers (Dynamic event.  Sunglint to
the right should have given good conditions [large disc] for documenting
the highly changeable river courses in the region.  ISS/CEO images
of three very remote rivers to the south [Beni and Mamore in Bolivia, and
Bermejo in N Argentina] show significant changes in pattern since the
Landsat base images used by analysts were acquired only 10-15 years
ago.  Two of these rivers have been written up, one for NASAís Earth
Observatory website this week [see below]),
Johnston Island
(fringing coral reefs around this remote island WSW of the
Tashkent, Uzbekistan (nadir pass), Cairo,
(nadir pass), Damascus, Syria (nadir pass),
Madrid, Spain (pointing a touch left), Barcelona,
(pointing a touch right), Tessina landslide site
(detailed images shot for ~30 secs under track should acquire this
mountainside research site),
Urumqui, China (nadir pass),
Berlin, Germany (opportunity to acquire this major city
London, England, Great Britain (clear skies,
despite rain to the north),
Missouri River sites (LEWIS
& CLARK SITES:  In this phase of attempting to get the detailed
L&C images, crew should try for a mapping swath of 2-3 frames looking
left up the Missouri River for several sites.  On May 14, 1804, the
explorers started up the Missouri river in a 55-ft keelboat and two
smaller craft),
Weeks Creek, California (this landslide
research site near San Francisco is under track:  which suggested a
mapping swath of 2-3 frames across the short distance from the Pacific
coast to San Francisco Bay), and
Seattle, Washington (nadir
CEO images can be viewed at the websites
See also the website "Space Station Challenge"

U.S. and Russian Segment Status (as of 1:55pm EST).

Environmental Control and Life Support (ECLSS) and Thermal Control
Elektron O2 generator is powered Off.  Vozdukh CO2 scrubber
is On (in Manual Mode 5/3).  U.S. CDRA CO2 scrubber is Off. 
TCCS (trace contaminant control subsystem) is operating.  MCA (major
constituents analyzer) is operating.  BMP Harmful Impurities unit:
absorbent bed #1 in Purify mode, bed #2 in Purify mode.  RS air
conditioner SKV-1 is On; SKV-2 is Off.

SM Working Compartment:  Pressure (mmHg) — 742; temperature
(deg C) — 26.4; ppO2 (mmHg) — data invalid; ppCO2 (mmHg) —
data invalid.
SM Transfer Compartment:  Pressure (mmHg) — 741; temperature
(deg C) — 20.4.
FGB Cabin:  Pressure (mmHg) — 744; temperature (deg C) —
Node:  Pressure (mmHg) — 741.34; temperature (deg C) — 23.7
(shell); ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
U.S. Lab:  Pressure (mmHg) — 743.1; temperature (deg C) —
23.0; ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
Joint Airlock (Equip. Lock):  Pressure (mmHg) — 743.3;
temperature (deg C) — 27.6; shell heater temp (deg C) — 24.2, ppO2
(mmHg) — 167.2; ppCO2 (mmHg) — 3.5.
PMA-1:  Shell heater temp (deg C) — 24.6
PMA-2:  Shell heater temp (deg C) — 14.3.
         (n/a = data not

Propulsion System (PS):
Total propellant load available: 3646 kg (8038 lb) as
of 7/31 
[SM(774) + FGB(2872) + Progress M(0) +Progress
M-1(0)].  (Capability: SM — 860 kg; FGB — 6120 kg).

Electrical Power Systems (EPS):
Both P6 channels fully operational.  BGA (beta gimbal
assembly) 2B and 4B both in Autotrack (solar tracking with 45 deg bias
for drag reduction).
SM batteries:  Battery #5 is off line(Capacity Restoration mode,
ROM); battery #7 is in "Cycle" mode; all other batteries (6)
are in "Partial Charge" mode.  Batteries #1and #3 showing
degraded capacity. 
FGB batteries:  Battery #1 is off line (Capacity Restoration mode,
ROM); all other batteries (5) are in "Partial Charge"
Plasma Contactor Unit PCU-1 in Standby mode; PCU-2 in Standby mode.

Command & Data Handling Systems:
C&C-1 MDM is prime, C&C-2 is back-up, and C&C-3 is in
GNC-1 MDM is prime; GNC-2 is Backup.
INT-1 is operating; INT-2 is Off.
EXT-2 is On (primary), EXT-1 is Off (both now upgraded to R3).
LA-1, LA-2 and LA-3 MDMs are all operating.
PL-2 MDM is Off; PL-1 MDM is Operational.
APS-1 (automated payload switch #1) and APS-2 are both On.
SM Terminal Computer (TVM): 3 redundant lanes (of 3) operational.
SM Central Computer (TsVM): 3 redundant lanes (of 3) operational.

Attitude Control Systems:
3 CMGs on-line (CMG-1 failed).
State vector source — U.S. SIGI-1 (GPS)
Attitude source — U.S. SIGI-1 (GPS)
Angular rate source — RGA-1

Flight Attitude:
XPOP (x-axis perpendicular to orbit plane = ìsun-fixedî [yaw: 0.5
deg, pitch: -9.0 deg., roll: 0 deg]), with CMG TA (thruster
Solar Beta angle:  25.6 deg (magnitude increasing).

Communications & Tracking Systems:
FGB MDM-1 is powered Off; FGB MDM-2 is operational.
All other Russian communications & tracking systems are
S-band is operating nominally (on string 2).
Ku-band is operating nominally.
Audio subsystem is operating nominally (IAC-2 is prime).
Video subsystem operating nominally.
HCOR (high-rate communications outage recorder) is operating

SSRMS/Canadarm2 based at MBS PDGF #1 with Keep Alive (KA) power
on both strings.
MBS: KA power on both strings. 
MT: latched and mated at WS4. 
POA: KA power on both strings.
RWS (robotics workstations): Lab RWS is On (DCP connected); Cupola RWS is

ISS Orbit  (as of this morning, 4:13am EDT [=
Mean altitude — 384.2 km
Apogee — 389.6 km
Perigee — 378.8 km
Period — 92.24 min.
Inclination (to Equator) —  51.63 deg
Eccentricity — 0.0008036
Orbits per 24-hr. day — 15.61
Solar Beta Angle — 25.6 deg (magnitude increasing)
Mean altitude loss in last 24 hours — 100 m
Revolutions since FGB/Zarya launch (Nov. í98)  — 26992
For more on ISS orbit and worldwide naked-eye visibility dates/times,

SpaceRef staff editor.